OpenStudy (anonymous):

1. 8y 7 ÷ 4y 5 2. (18x 5 + 6x 4 - 12x 3) ÷ 6x 2 helppppp???

OpenStudy (anonymous):

\[\frac{x^n}{x^m}=x^{n-m}\]

OpenStudy (anonymous):

huh ?

OpenStudy (anonymous):

so first take care of the constants, \[\frac{8}{4}=\]

OpenStudy (anonymous):

2

OpenStudy (anonymous):

then you have \[\frac{y^7}{y^5}\] since they have the same base you can use the above property so \[\frac{y^7}{y^5}=y^{7-5}\]

OpenStudy (anonymous):

so combine your constant and your variables to get \[2y^{7-5}=2y^2\]

OpenStudy (anonymous):

\[ (18x ^{5}+6x ^{4}-12x ^{3})\div6x ^{2} \] is the q btw it didnt post the powers

OpenStudy (anonymous):

the second equation you need to split everything up into their own fractions

OpenStudy (anonymous):

same for the first one

OpenStudy (anonymous):

I have no idea how to do this

OpenStudy (anonymous):

so \[\frac{18x^5+6x^4-12x^3}{6x^2}=\frac{18x^5}{6x^2}+\frac{6x^4}{6x^2}-\frac{12x^3}{6x^2}\]

OpenStudy (anonymous):

then do what we did in the first one

OpenStudy (anonymous):

Still have no idea lol

OpenStudy (anonymous):

Thanks guys